OpenStudy (anonymous):

7a+3b=85si a=2b aflati numerele naturale a si b

OpenStudy (angela210793):

I think this is what you mean..... 7*2b+3b=85 14b+3b=85 17b=85 b=85/17

OpenStudy (anonymous):

multumesc

OpenStudy (radar):

7a+3b=85 a=2b 14b+3b=85 17b=85 b=5 a=2b=6